OpenStudy (anonymous):

6x^4 - 18x^2 +6 = 0 Solve for 0.

OpenStudy (tanya123):

Simplifying 6x4 + -18x2 + 6 = 0 Reorder the terms: 6 + -18x2 + 6x4 = 0 Solving 6 + -18x2 + 6x4 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '6'. 6(1 + -3x2 + x4) = 0 Ignore the factor 6

OpenStudy (anonymous):

How do I solve x^4-3x^2+1?

OpenStudy (anonymous):

@tanya123

OpenStudy (thomas5267):

Let \(u=x^2\). Rewrite the equation in terms of \(u\).

OpenStudy (anonymous):

a^2 - 3a+1

OpenStudy (anonymous):

or u^2 - 3u +1

OpenStudy (thomas5267):

Use the quadratic equation.

OpenStudy (anonymous):

ooo right

OpenStudy (anonymous):

Nothing happen to the u's being back to x^2? I checked the solutions for 6x^4-18x^2+6=0 and I get 4 solutions.

OpenStudy (anonymous):

@thomas5267

OpenStudy (anonymous):

but when I solve for u I get only 2 solutions

OpenStudy (thomas5267):

Note that if \(x^2=n\), then \(x=\sqrt{n}\) and \(x=-\sqrt{n}\).

OpenStudy (anonymous):

I got confused by that

OpenStudy (thomas5267):

Consider this case. If \(x^2=9\), clearly \(x=3\) is a solution. However, \(x=-3\) is also a solution as \((-3)^2=9\).

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

but in the case after getting the solutions of u after using the quadratic formula. What do I do to get the other 2?

OpenStudy (anonymous):

I got x= 1/2(3-sqrt(5)) and x=1/2(3+sqrt(5))

OpenStudy (anonymous):

after solving for u

OpenStudy (anonymous):

but the real solutions that I need from 6x^4 -18x^2 +6 = 0 are 4

OpenStudy (thomas5267):

\[ 6(x^4-3x^2+1)=0\\ u=x^2\\ \begin{align*} u^2-3u+1&=0\\ u&=\frac{3\pm\sqrt{(-3)^2-4}}{2}\\ u&=\frac{3\pm\sqrt{5}}{2}\\ u&=\frac{3+\sqrt{5}}{2} \text{ or } u=\frac{3-\sqrt{5}}{2} \end{align*} \]

OpenStudy (anonymous):

ooo

OpenStudy (anonymous):

the u's become x^2 and then I get 2 more solutions. But how do I simplify the square root over the solutions of u?

OpenStudy (thomas5267):

\[ \begin{align*} u&=\frac{3+\sqrt{5}}{2}&&\\ x^2&=\frac{3+\sqrt{5}}{2}&\text{ or } x^2&=\frac{\sqrt{5}-3}{2}\\ x&=\sqrt{\frac{3+\sqrt{5}}{2}}&\text{ or } x&=\sqrt{\frac{\sqrt{5}-3}{2}}\\ \end{align*} \] The problem now is how to simplify the square root in square root.

OpenStudy (anonymous):

yes. I got that. I'm trying to simplify that now but I don't know how. I haven't been taught how to do that

OpenStudy (anonymous):

\[x=\frac{ 1+\sqrt{5} }{ 2 }\]

OpenStudy (anonymous):

How can I simplify it to that?

OpenStudy (anonymous):

I'm aware there are 3 more

OpenStudy (thomas5267):

I found this document on the internet. This document is about how to solve nested square roots. http://www.cybertester.com/data/denest.pdf

OpenStudy (thomas5267):

@ganeshie8 How to solve nested square roots? I need a coffee break now lol.

OpenStudy (thomas5267):

Theorem (4.12) in that document might be helpful.

OpenStudy (thomas5267):

Formula given in the link: \[ \sqrt{A\pm B}=\sqrt{\frac{1}{2}A+\frac{1}{2}\sqrt{A^2-B^2}}\pm\sqrt{\frac{1}{2}A-\frac{1}{2}\sqrt{A^2-B^2}} \text{, given } A,B\in\mathbb{R} \\\text{ and }A>B>0 \] \[ \begin{align*} x=\sqrt{\frac{3}{2}+\frac{\sqrt{5}}{2}}&=\sqrt{\frac{3}{4}+\frac{1}{2}\sqrt{\left(\frac{3}{2}\right)^2-\left(\frac{\sqrt{5}}{2}^2\right)}}+\sqrt{\frac{3}{4}-\frac{1}{2}\sqrt{\left(\frac{3}{2}\right)^2-\left(\frac{\sqrt{5}}{2}^2\right)}}\\ &=\sqrt{\frac{3}{4}+\frac{1}{2}\sqrt{\frac{9}{4}-\frac{5}{4}}}+\sqrt{\frac{3}{4}-\frac{1}{2}\sqrt{\frac{9}{4}-\frac{5}{4}}}\\ &=\sqrt{\frac{3}{4}+\frac{1}{2}}+\sqrt{\frac{3}{4}-\frac{1}{2}}\\ &=\sqrt{\frac{5}{4}}+\sqrt{\frac{1}{4}}\\ &=\frac{\sqrt{5}}{2}+\frac{1}{2}\\ &=\frac{1+\sqrt{5}}{2} \end{align*} \]

OpenStudy (anonymous):

<3

OpenStudy (thomas5267):

In case you can't read the overflowed part. \[ \begin{align*} x&=\sqrt{\frac{3}{2}+\frac{\sqrt{5}}{2}}\\ &=\sqrt{\frac{3}{4}+\frac{1}{2}\sqrt{\left(\frac{3}{2}\right)^2-\left(\frac{\sqrt{5}}{2}^2\right)}}+\sqrt{\frac{3}{4}-\frac{1}{2}\sqrt{\left(\frac{3}{2}\right)^2-\left(\frac{\sqrt{5}}{2}^2\right)}}\\ &=\sqrt{\frac{3}{4}+\frac{1}{2}\sqrt{\frac{9}{4}-\frac{5}{4}}}+\sqrt{\frac{3}{4}-\frac{1}{2}\sqrt{\frac{9}{4}-\frac{5}{4}}}\\ &=\sqrt{\frac{3}{4}+\frac{1}{2}}+\sqrt{\frac{3}{4}-\frac{1}{2}}\\ &=\sqrt{\frac{5}{4}}+\sqrt{\frac{1}{4}}\\ &=\frac{\sqrt{5}}{2}+\frac{1}{2}\\ &=\frac{1+\sqrt{5}}{2} \end{align*} \]

OpenStudy (thomas5267):

I am probably overcomplicating things as usual, @ganeshie8 . It is absolutely common for me to do so. Any other solutions?

OpenStudy (thomas5267):

You now have to do this to all four solutions... Good luck... :( I have to take a coffee break.

OpenStudy (anonymous):

Yeah, I got it. I am falling sleep but still work to do haha

OpenStudy (anonymous):

Thanks you very much. After this I will have an even more problematic one. Idk if anyone would be able to help me out on that one.

OpenStudy (anonymous):

@thomas5267 I have a question regarding this again. When I solve for sqrt(3-sqrt(5)/2) wht changes on the equation?

OpenStudy (anonymous):

the above one is 3/2 + sqrt5/2 but now i have 3/2 - sqrt5/2